# TabSTAR: A Foundation Tabular Model With Semantically Target-Aware Representations

### v1

@@ E.2.2 CARTE
E.2.3 CatBoost and CatBoost-Tuned
We run CatBoost using the catboost package [^33] https://pypi.org/project/catboost/ and run the default configuration suggested by [25] by setting $early\_stopping\_rounds=50$, $od\_pval=0.001$, $iterations=2000$. For the tuned version, we use the Optuna package [^34] https://pypi.org/project/optuna/ with random search, with a budget of 4 hours for every run and parallelizing trials on 8 CPU cores. We use 5-fold cross-validation and take the best configuration selected based on this mean score. We use it then to retrain the model on the full training data. For hyperparameter space, we follow the hyperparameter search suggested by [37] as detailed in Table 15.
@@ E.2.3 CatBoost and CatBoost-Tuned
E.2.4 XGBoost and XGBoost-Tuned
We run XGBoost using the xgboost package [^35] https://pypi.org/project/xgboost/ and follow the same procedure as for CatBoost, except additional preprocessing (e.g., transforming categorical variables into one-hot encodings). For the default configuration, we follow the suggestion of [25] and use: $booster="gbtree"$, $early\_stopping\_rounds=50$, $n\_estimators=2000$. For the tuned variant, we follow the hyperparameter search space suggested by [37], as shown in Table 16.
@@ E.2.4 XGBoost and XGBoost-Tuned
E.2.5 Random Forest
We treat Random Forest as a weak baseline to establish a lower-bound reference for each dataset split. We run it with the sklearn package  [^36] https://scikit-learn.org/ and use its default configuration with $n\_estimators=100$.
@@ Appendix F Extended Main Results
In this section, we often use abbreviated model names for conciseness: We refer to CatBoost variants as CatB and CatB-T, with the latter being the Tuned version. Similarly, we use XGB and XGB-T for XGBoost. We maintain the abbreviation of Random Forest (RF) and shorten TabPFN-v2 to simply TabPFN. For models that run on an unlimited number of training examples, we add a -U suffix, i.e. TabSTAR-U, CatB-T-U, and XGB-T-U.
F.1 Dataset Level Performance
We report AUROC for classification and $R^{2}$ for regression, with 95% CIs computed over the 10 runs for each dataset. Tables  17 and 18 summarize classification performance on datasets with up to 10K and over 10K examples, respectively. Tables 19 and 20 to the same for regression tasks. For conciseness, datasets are referred by their ID from Appendix D.

### v2

@@ E.2.4 CARTE
E.2.5 RealMLP
We run RealMLP using its official implementation in the pytabkit package. [^36] https://pypi.org/project/pytabkit/ Following the advice of its authors, we disable label smoothing and optimize for cross_entropy for binary classification and 1-auc_ovr for multiclass classification, and keep all other default hyperparameters. For the tuned version, we follow the hyperparameter search space suggested by [39] as detailed in Table 15.
Table 15: RealMLP-Tuned Hyperparameters Search Space
Hyperparameter Search Space
num_emb_type Choice([None, PBLD, PL, PLR])
add_front_scale Choice([True, False], p=[0.6, 0.4])
lr $\log\mathcal{U}(2e-2,3e-1)$
p_drop Choice([0.0, 0.15, 0.3], p=[0.3, 0.5, 0.2])
act Choice([ReLU, SELU, Mish])
hidden_sizes Choice([[256, 256, 256], [64, 64, 64, 64], [512]], p=[0.6, 0.2, 0.2])
wd Choice([0.0, 2e-2])
plr_sigma $\log\mathcal{U}(0.05,0.5)$ if classification else 0.1
E.2.6 CatBoost
We run CatBoost using the catboost package [^37] https://pypi.org/project/catboost/ and run the default configuration suggested by [26] by setting early_stopping_rounds=50early\_stopping\_rounds=50, od_pval=0.001od\_pval=0.001, $iterations=2000$. For the tuned version, we follow the hyperparameter search suggested by [38] as detailed in Table 16.
@@ E.2.6 CatBoost
E.2.7 XGBoost
We run XGBoost using the xgboost package. [^38] https://pypi.org/project/xgboost/ For the default configuration, we follow the suggestion of [26] and use: booster=gbtreebooster="gbtree", early_stopping_rounds=50early\_stopping\_rounds=50, n_estimators=2000n\_estimators=2000. For the tuned variant, we follow the hyperparameter search space suggested by [38], as shown in Table 17.
@@ E.2.7 XGBoost
E.2.8 LightGBM
We run LightGBM using the lightgbm package. [^39] https://pypi.org/project/lightgbm/ and its default implementation. For the tuned variant, we follow the hyperparameter search space suggested by [38], as shown in Table 18.
Table 18: LightGBM-Tuned Hyperparameters Search Space
Hyperparameter Search Space
num_leaves $\mathcal{U}\{5,6,\ldots,50\}$
max_depth $\mathcal{U}\{3,4,\ldots,20\}$
learning_rate $\log\mathcal{U}(e^{-3},1)$
n_estimators $\mathcal{U}\{50,51,\ldots,2000\}$
min_child_weight $\{10^{-5},10^{-3},10^{-2},10^{-1},1,10,10^{2},10^{3},10^{4}\}$
subsample $\mathcal{U}(0.2,0.8)$
colsample_bytree $\mathcal{U}(0.2,0.8)$
reg_alpha $\{0,10^{-1},1,2,5,7,10,50,100\}$
reg_lambda $\{0,10^{-1},1,5,10,20,50,100\}$
E.2.9 Random Forest
We treat Random Forest as a weak baseline to establish a lower-bound reference for each dataset split. We run it with the sklearn package  [^40] https://scikit-learn.org/ and use its default configuration with n_estimators=100n\_estimators=100.
@@ Appendix F Extended Main Results
F.1 Technical Limitations